# Multiview learning with twin parametric margin SVM

### v1

@@ 3.2 Nonlinear MvTPMSVM
The dual formulations and solutions of the problem (3.2) and (3.2) can be calculated in a similar way as in the linear case, and are given by:
$\underset{\tilde{\tau}}{\min}\hskip 2.84544pt\leavevmode\nobreak$ $\frac{1}{2}\tilde{\tau}^{T}\tilde{\Lambda}\tilde{\tau}+\kappa_{1}\tilde{\tau}$
@@
$\underset{\hat{\tau}}{\min}\hskip 2.84544pt\leavevmode\nobreak$ $\frac{1}{2}\hat{\tau}^{T}\hat{\Lambda}\hat{\tau}+\kappa_{2}\hat{\tau}$
@@ 4.1 Experimental setup
The experimental hardware setup comprises a PC with an Intel(R) Xeon(R) Gold 6226R CPU running at $2.90$GHz and equipped with $128$ GB of RAM, running Windows $11$ operating system possessing Matlab R2023a. The dataset is randomly partitioned into a ratio of $70:30$, allocating $70\%$ of the data for training and $30\%$ for testing. We employ a five-fold cross-validation technique and a grid search approach to optimize the hyperparameters of the models. For all experiments, we opt the Gaussian kernel function represented by $k(x_{i},x_{j})=e^{-\frac{\|x_{i}-x_{j}\|}{2\sigma^{2}}}$ for each model. The kernel parameter $\sigma$ is selected from the following range: $\{2^{-5},2^{-4},\ldots 2^{4},2^{5}\}$. We adopt equal penalty parameters, i.e., $C_{1}=C_{3}$ and $C_{2}=C_{4}=D_{1}=D_{2}$, to mitigate computational costs and are selected from the range $\{2^{-5},2^{-4},\ldots 2^{4},2^{5}\}$. For the baseline MvTSVMs and MVNPSVM model, we set $C_{1}=C_{2}=C_{3}=C_{4}$ and selected from the range $\{2^{-5},2^{-4},\ldots 2^{4},2^{5}\}$. In SVM-2K and PSVM-2V, we set $C_{1}=C_{1}=D$ and selected from the range $\{2^{-5},2^{-4},\ldots 2^{4},2^{5}\}$. For MVLDM, the parameter $C_{1},v_{1},v_{2},\theta,\sigma$ are chosen from $\{2^{-5},2^{-4},\ldots,2^{4},2^{5}\}$. The parameter $\epsilon$ in the proposed MvTPMSVM model along with the baseline models is set to $0.1$. The generalization performance of the proposed MvTPMSVM model has been evaluated by comparing it with baseline models across various metrics including $accuracy$, $sensitivity$, $precision$, $specificity$, and $error$ $rate$. Mathematically,
$Accuracy\leavevmode\nobreak\ (ACC)=\frac{\mathcal{TN}+\mathcal{TP}}{\mathcal{FP}+\mathcal{FN}+\mathcal{TP}+\mathcal{TN}},$ (55)
$Sensitivity\leavevmode\nobreak\ (Seny)=\frac{\mathcal{TP}}{\mathcal{FN}+\mathcal{TP}},$ (56)
$Precision\leavevmode\nobreak\ (Pren)=\frac{\mathcal{TP}}{\mathcal{FP}+\mathcal{TP}},$ (57)
$Specificity\leavevmode\nobreak\ (Spey)=\frac{\mathcal{TN}}{\mathcal{TN}+\mathcal{FP}},$ (58)
@@
where true positive ($\mathcal{TP}$) represents the count of patterns belonging to positive class that are accurately classified, while false negative ($\mathcal{FN}$) signifies the count of patterns belonging to positive class that are inaccurately classified, false positive ($\mathcal{FP}$) denotes the count of patterns belonging to negative class that are inaccurately classified, and true negative ($\mathcal{TN}$) describes the number of data points of negative class that are correctly classified.
@@ 4.2 Experiments on synthetic datasets
Table 2: Classification performance of proposed MvTPMSVM along with the baseline models based on classification accuracy using non-linear kernel for synthetic datasets.
@@
Table 2 displays the comparative experimental outcomes of the proposed MvTPMSVM model, along with the baseline models encompassing non-linear case. From Table 2, the average accuracy of the proposed MvTPMSVM model is $97.78\%$, surpassing the performance of the baseline models. It is evident that the proposed MvTPMSVM model consistently outperforms the baseline models. Thus, the experimental comparison of datasets demonstrates the superior performance of our proposed MvTPMSVM model.
@@ 4.3 Experiments on real-world UCI and KEEL datasets
Table 3 shows the experimental results including ACC, Seny, Spey, and Pren of the proposed MvTPMSVM model along with the baseline SVM-2K, MvTSVM, MVNPSVM, and PSVM-2V models. The optimal parameters of the proposed MvTPSVM model and the baseline models corresponding to the error rates are presented in Table 5. The comparison in terms of ACC, Seny, Spey, Pren and error rates indicate that our proposed MvTPMSVM model yields better performance than the baseline models on most of the datasets. From Table 3, the average ACC of the proposed MvTPMSVM model is $85.55\%$. In contrast, the average ACC for SVM-2K, MvTSVM, MVNPSVM, PSVM-2V and MVLDM models stands at $84.48\%$, $82.25\%$, $80.89\%$, $83.22\%$ and $84.32\%$, respectively. As the average ACC can be influenced by exceptional performance in one dataset that compensates for losses across multiple datasets, it might be considered a biased measure. Hence, we utilize the ranking method to assess the effectiveness and evaluate the performance of the models. Here, each classifier is assigned a rank, where the model exhibiting superior performance receives a lower rank, and the model with inferior performance is assigned a higher rank. In the evaluation of $k$ models across $N$ datasets, $r_{j}^{i}$ represents the rank of the $j^{th}$ model on the $i^{th}$ dataset. $R_{j}=\frac{1}{N}\sum_{i=1}^{N}r_{j}^{i}$ calculates the average rank of $j^{th}$ model. The average ranks for SVM-2K, MvTSVM, MVNPSVM, PSVM-2V, MVLDM and MvTPMSVM are $3.25$, $4.29$, $4.62$, $3.89$, $2.87$ and $2.07$, respectively. The evident observation is that the proposed MvTPMSVM model demonstrates the most favorable average ranks. Hence, the generalization ability of the proposed MvTPMSVM model surpasses that of the baseline models. Now we conduct the statistical tests to determine the significance of the results. To evaluate if there are statistically significant differences between the models, we employ the Friedman test [3]. The Friedman test involves comparing the average ranks attributed to different models, subsequently determining the presence of statistical distinctions among the models based on their respective rankings. The null hypothesis posits that all the models share an equal average rank, indicating comparable performance. The Friedman test adheres to the chi-squared distribution $\chi_{F}^{2}$ with $(k-1)$ degrees of freedom (d.o.f.), and its calculation involves: $\chi_{F}^{2}=\frac{12N}{k(k+1)}\left[\sum_{j}R_{j}^{2}-\frac{k(k+1)^{2}}{4}\right]$. The Friedman statistic $(F_{F})$ is calculated as: $F_{F}=\frac{(N-1)\chi_{f}^{2}}{N(k-1)-\chi_{F}^{2}}$, where the d.o.f. for the $F$-distribution is $(k-1)$ and $(N-1)\times(k-1)$. At the $5\%$ level of significance, we obtain $\chi_{F}^{2}=70.1627$ and $F_{F}=18.4965$ for $k=6$ and $N=55$. With reference to $F_{F}(5,270)=2.4134$ from statistical $F$-distribution table. We reject the null hypothesis as $18.4965>2.2474$. Hence, statistical difference exists among the compared models. Consequently, we employ the Nemenyi post hoc test to analyze the pairwise distinctions between the models. The critical difference $(C.D.)$ is computed as follows: $C.D.=q_{\alpha}\sqrt{\frac{k(k+1)}{6N}}$, where $q_{\alpha}$ represents the critical value from the distribution table for the two-tailed Nemenyi test. Referring to the statistical $F$-distribution table, with $q_{\alpha}=2.850$ at a $5\%$ significance level, the calculated $C.D.$ is $1.0167$. The differences in average ranks between the proposed MvTPMSVM model and the baseline SVM-2K, MvTSVM, MVNPSVM, PSVM-2V, and MVLDM models are $1.18$, $2.22$, $2.55$, $1.82$ and $0.80$ respectively. The average rank difference between the proposed MvTPMSVM is greater than the ranks of the SVM-2K, MvTSVM, MVNPSVM, and PSVM-2V models. According to the Nemenyi post hoc test, noteworthy differences are observed between the proposed MvTPMSVM model and the baseline models (except MVLDM). The MvTPMSVM model surpasses MVLDM in terms of average rank. The superior performance of the proposed MvTPMSVM model in comparison to the baseline models is evident.
@@
Table 5: Error rate and optimal parameters of the proposed MvTPMSVM model along with the baseline models on UCI and KEEL datasets with the non-linear kernel.
@@
Dataset $\downarrow$ Error rate Error rate Error rate Error rate Error rate Error rate
@@
Dataset $\downarrow$ Error rate Error rate Error rate Error rate Error rate Error rate
@@
Average Error Rate $0.1678$ $0.1775$ $0.1678$ $0.1911$ $0.1552$ $0.1445$
@@
Furthermore, in the analysis of models, we employ the pairwise win-tie-loss (W-T-L) sign test. According to the W-T-L sign test, the null hypothesis posits that two models perform equivalently and are expected to secure victories in $N/2$ datasets, where $N$ denotes the count of datasets. If the classification model secures wins in approximately $N/2+1.96\sqrt{N}/2$ datasets, then the model is deemed significantly better. In the event of an even number of ties between two models, these ties are evenly distributed between them. Nevertheless, if the number of ties is odd, one tie is disregarded, and the remaining ties are distributed among the specified classifiers. In this instance, with $N=55$, the occurrence of wins in at least $42.035$ datasets by one of the models signifies a noteworthy distinction between the models. Table 7 presents the outcomes in terms of pairwise wins, ties, and losses across UCI and KEEL datasets. In each entry, $[x,y,z]$, $x$ denotes the number of times the model specified in the corresponding row secures a win, $y$ represents the number of ties, and $z$ signifies the occurrences of losses in comparison to the model mentioned in the respective column. Upon meticulous examination of Table 7, it is apparent that the proposed MvTPMSVM model outperforms the baseline MvTSVM, MVNPSVM, and PSVM-2V models by achieving a respective number of victories: $47$, $47$, and $43$, out of a total of $55$ datasets. The proposed MvTPMSVM demonstrates a statistically significant distinction compared to the baseline models, except for SVM-2K and MVLDM. Exhibiting a notable level of performance, the MvTPMSVM model succeeds in $39$ out of $55$ datasets over SVM-2K and $26$ out of $55$ datasets over MVLDM. However, the winning percentage of the MvTPMSVM model demonstrates its effectiveness in comparison to the SVM-2K and MVLDM models. The evidence demonstrates that the proposed MvTPMSVM model exhibits significant superiority when compared to the baseline models.
@@ 4.4 Experiments on real-world AwA datasets
Table 8 displays the ACC, Seny, Spey, and Pren of the proposed MvTPMSVM model with SVM-2K, MvTSVM, MVNPSVM, PSVM-2V, and MVLDM. The optimal parameters of the proposed MvTPSVM model and the baseline models corresponding to the error rates are presented in Table 10. From Table 8, it is evident that the proposed MvTPMSVM model exhibits superior generalization performance across the majority of datasets. The average ACC of the proposed MvTPMSVM model and the baseline SVM-2K, MvTSVM, MVNPSVM, PSVM-2V, and MVLDM models are $73.57\%$, $71.02\%$, $69.4\%$, $70.50\%$, $71.02\%$ and $73.33\%$, respectively. Table 8 presents the average ranks of all models, determined by their accuracy values. It is worth noting that among all the models, our proposed MvTPMSVM model has the lowest average rank. Moreover, we perform the Friedman statistical test, followed by Nemenyi post hoc tests. For the significance level of $\alpha=0.05$, we calculate $\chi_{F}^{2}=50.81$, $F_{F}=12.8345$, and $F_{F}(5,220)=2.2550$. The null hypothesis is rejected as $F_{F}(5,220)<F_{F}$. Now, the Nemenyi post-hoc test is employed to identify significant differences among the pairwise comparisons. We calculate $C.D.=1.1240$, indicating that the average rankings of the models in Table 8 should have a minimum difference by $1.1240$ to be considered statistically significant. The differences in average ranks between the proposed MvTPMSVM model and the baseline SVM-2K, MvTSVM, MVNPSVM, PSVM-2V, and MVLDM models are $1.60$, $2.85$, $2.26$, $1.89$ and $0.81$, respectively. $C.D.$ is exceeded by the observed differences. Thus, as per the Nemenyi post hoc test, the proposed MvTPMSVM model, exhibits significant distinctions from SVM-2K, MvTSVM, MVNPSVM, and PSVM-2V except MVLDM. Consequently, the proposed MvTPMSVM model demonstrates superior performance compared to baseline models except MVLDM. However, the proposed MvTPMSVM model surpasses MVLDM in terms of average rank. The evident superiority of the proposed MvTPMSVM model is evident when contrasted with the baseline models.
Table 10: Error rate and optimal parameters of the proposed MvTPMSVM model along with the baseline models on AwA datasets with the non-linear kernel.
@@
Dataset $\downarrow$ Error rate Error rate Error rate Error rate Error rate Error rate
@@
Dataset $\downarrow$ Error rate Error rate Error rate Error rate Error rate Error rate
@@
Average Error Rate $0.2808$ $0.3060$ $0.2898$ $0.2950$ $0.2667$ $0.2643$
@@
Subsequently, additional experiments were performed by adjusting the sizes of the training sets, and their respective performance is illustrated in Figures 3. We employ a random selection method, using $30\%$ of the samples for testing, while systematically altering the set sizes allocated for training. Figure 3, exhibiting different training sizes, consistently demonstrates that the proposed MvTPMSVM model generally outperforms other baseline models. Moreover, as the training sizes increase, the MvTPMSVM model exhibits superior performance compared to baseline models.
Based on Figure 5, the Seny, Spey, and Pren analysis indicates that the performance of the proposed MvTPMSVM model achieved the highest position in the baseline models, demonstrating its competitive nature. The average Seny, Spey, and Pren are $79.65$, $78.34$, and $80.36$, respectively. The statistical tests provide further evidence that the proposed models outperformed all other models in the AwA datasets and achieved the highest rankings in Seny, Spey, and Pren analyses.

### v2

@@ 3.2 Nonlinear MvTPMSVM
The dual formulations of (3.2) and (3.2) can be calculated in a similar way as in the linear case, and are given by:
$\underset{\tilde{\tau}}{\max}\hskip 2.84544pt\leavevmode\nobreak$ $\frac{1}{2}\tilde{\tau}^{T}\tilde{\Lambda}\tilde{\tau}+\kappa_{1}\tilde{\tau}$
@@
$\underset{\hat{\tau}}{\max}\hskip 2.84544pt\leavevmode\nobreak$ $\frac{1}{2}\hat{\tau}^{T}\hat{\Lambda}\hat{\tau}+\kappa_{2}\hat{\tau}$
@@ 4.1 Experimental setup
The experimental hardware setup comprises a PC with an Intel(R) Xeon(R) Gold 6226R CPU running at $2.90$GHz and equipped with $128$ GB of RAM, running Windows $11$ operating system possessing Matlab R2023a. The dataset is randomly partitioned into a ratio of $70:30$, allocating $70\%$ of the data for training and $30\%$ for testing. We employ a five-fold cross-validation technique and a grid search approach to optimize the hyperparameters of the models. For all experiments, we opt the Gaussian kernel function represented by $k(x_{i},x_{j})=e^{-\frac{\|x_{i}-x_{j}\|}{2\sigma^{2}}}$ for each model. The kernel parameter $\sigma$ is selected from the following range: $\{2^{-5},2^{-4},\ldots 2^{4},2^{5}\}$. We adopt equal penalty parameters, i.e., $C_{1}=C_{3}$ and $C_{2}=C_{4}=D_{1}=D_{2}$, to mitigate computational costs and are selected from the range $\{2^{-5},2^{-4},\ldots 2^{4},2^{5}\}$. For the baseline MvTSVM and MVNPSVM model, we set $C_{1}=C_{2}=C_{3}=C_{4}$ and selected from the range $\{2^{-5},2^{-4},\ldots 2^{4},2^{5}\}$. In SVM-2K and PSVM-2V, we set $C_{1}=C_{1}=D$ and selected from the range $\{2^{-5},2^{-4},\ldots 2^{4},2^{5}\}$. For MVLDM, the parameter $C_{1},v_{1},v_{2},\theta,\sigma$ are chosen from $\{2^{-5},2^{-4},\ldots,2^{4},2^{5}\}$. The parameter $\epsilon$ in the proposed MvTPMSVM model along with the baseline models is set to $0.1$. The generalization performance of the proposed MvTPMSVM model has been evaluated by comparing it with baseline models across various metrics including $accuracy$, $sensitivity$, $precision$, $specificity$, and $error$ $rate$. Mathematically,
$Accuracy\leavevmode\nobreak\ (\text{ACC})=\frac{\mathcal{TN}+\mathcal{TP}}{\mathcal{FP}+\mathcal{FN}+\mathcal{TP}+\mathcal{TN}},$ (55)
$Sensitivity\leavevmode\nobreak\ (\text{Seny})=\frac{\mathcal{TP}}{\mathcal{FN}+\mathcal{TP}},$ (56)
$Precision\leavevmode\nobreak\ (\text{Pren})=\frac{\mathcal{TP}}{\mathcal{FP}+\mathcal{TP}},$ (57)
$Specificity\leavevmode\nobreak\ (\text{Spey})=\frac{\mathcal{TN}}{\mathcal{TN}+\mathcal{FP}},$ (58)
@@
where true positive ($\mathcal{TP}$) represents the count of patterns belonging to $+1$ class that are accurately classified, while false negative ($\mathcal{FN}$) signifies the count of patterns belonging to $-1$ class that are inaccurately classified, false positive ($\mathcal{FP}$) denotes the count of patterns belonging to $-1$ class that are inaccurately classified, and true negative ($\mathcal{TN}$) describes the number of data points of $-1$ class that are correctly classified.
@@ 4.2 Experiments on synthetic datasets
Table 2: Performance comparison of the proposed MvTPMSVM along with the baseline models using non-linear kernel for synthetic datasets.
@@
Table 2 displays the comparative experimental outcomes of the proposed MvTPMSVM model, along with the baseline models encompassing non-linear case. From Table 2, the average ACC of the proposed MvTPMSVM model is $97.78\%$, surpassing the performance of the baseline models. It is evident that the proposed MvTPMSVM model consistently outperforms the baseline models. Thus, the experimental comparison of datasets demonstrates the superior performance of our proposed MvTPMSVM model.
@@ 4.3 Experiments on real-world UCI and KEEL datasets
Table 3 shows the experimental results including ACC, Seny, Spey, and Pren of the proposed MvTPMSVM model along with the baseline SVM-2K, MvTSVM, MVNPSVM, and PSVM-2V models. The optimal parameters of the proposed MvTPSVM model and the baseline models corresponding to the $error\leavevmode\nobreak\ rates$ are presented in Table 5. The comparison in terms of ACC, Seny, Spey, Pren and $error\leavevmode\nobreak\ rates$ indicate that our proposed MvTPMSVM model yields better performance than the baseline models on most of the datasets. From Table 3, the average ACC of the proposed MvTPMSVM model is $85.55\%$. In contrast, the average ACC for SVM-2K, MvTSVM, MVNPSVM, PSVM-2V and MVLDM models stands at $84.48\%$, $82.25\%$, $80.89\%$, $83.22\%$ and $84.32\%$, respectively. As the average ACC can be influenced by exceptional performance in one dataset that compensates for losses across multiple datasets, it might be considered a biased measure. Hence, we utilize the ranking method to assess the effectiveness and evaluate the performance of the models. Here, each classifier is assigned a rank, where the model exhibiting superior performance receives a lower rank, and the model with inferior performance is assigned a higher rank. In the evaluation of $k$ models across $N$ datasets, $r_{j}^{i}$ represents the rank of the $j^{th}$ model on the $i^{th}$ dataset. $R_{j}=\frac{1}{N}\sum_{i=1}^{N}r_{j}^{i}$ calculates the average rank of $j^{th}$ model. The average ranks for SVM-2K, MvTSVM, MVNPSVM, PSVM-2V, MVLDM and MvTPMSVM are $3.25$, $4.29$, $4.62$, $3.89$, $2.87$ and $2.07$, respectively. The evident observation is that the proposed MvTPMSVM model demonstrates the most favorable average ranks. Hence, the generalization ability of the proposed MvTPMSVM model surpasses that of the baseline models. Now we conduct the statistical tests to determine the significance of the results. To evaluate if there are statistically significant differences between the models, we employ the Friedman test [3]. The Friedman test involves comparing the average ranks attributed to different models, subsequently determining the presence of statistical distinctions among the models based on their respective rankings. The null hypothesis posits that all the models share an equal average rank, indicating comparable performance. The Friedman test adheres to the chi-squared distribution $\chi_{F}^{2}$ with $(k-1)$ degrees of freedom (d.o.f.), and its calculation involves: $\chi_{F}^{2}=\frac{12N}{k(k+1)}\left[\sum_{j}R_{j}^{2}-\frac{k(k+1)^{2}}{4}\right]$. The Friedman statistic $(F_{F})$ is calculated as: $F_{F}=\frac{(N-1)\chi_{f}^{2}}{N(k-1)-\chi_{F}^{2}}$, where the d.o.f. for the $F$-distribution is $(k-1)$ and $(N-1)\times(k-1)$. At the $5\%$ level of significance, we obtain $\chi_{F}^{2}=70.1627$ and $F_{F}=18.4965$ for $k=6$ and $N=55$. With reference to $F_{F}(5,270)=2.4134$ from statistical $F$-distribution table. We reject the null hypothesis as $18.4965>2.2474$. Hence, statistical difference exists among the compared models. Consequently, we employ the Nemenyi post hoc test to analyze the pairwise distinctions between the models. The critical difference $(C.D.)$ is computed as follows: $C.D.=q_{\alpha}\sqrt{\frac{k(k+1)}{6N}}$, where $q_{\alpha}$ represents the critical value from the distribution table for the two-tailed Nemenyi test. Referring to the statistical $F$-distribution table, with $q_{\alpha}=2.850$ at a $5\%$ significance level, the calculated $C.D.$ is $1.0167$. The differences in average ranks between the proposed MvTPMSVM model and the baseline SVM-2K, MvTSVM, MVNPSVM, PSVM-2V, and MVLDM models are $1.18$, $2.22$, $2.55$, $1.82$ and $0.80$ respectively. The average rank difference between the proposed MvTPMSVM with the baseline SVM-2K, MvTSVM, MVNPSVM, and PSVM-2V models are greater than the $C.D.$. According to the Nemenyi post hoc test, noteworthy differences are observed between the proposed MvTPMSVM model and the baseline models (except MVLDM). The MvTPMSVM model surpasses MVLDM in terms of average rank. The superior performance of the proposed MvTPMSVM model compared to the baseline models is evident.
@@
Table 5: $Error\leavevmode\nobreak\ rate$ and optimal parameters of the proposed MvTPMSVM model along with the baseline models on UCI and KEEL datasets with the non-linear kernel.
@@
Dataset $\downarrow$ $Error\leavevmode\nobreak\ rate$ $Error\leavevmode\nobreak\ rate$ $Error\leavevmode\nobreak\ rate$ $Error\leavevmode\nobreak\ rate$ $Error\leavevmode\nobreak\ rate$ $Error\leavevmode\nobreak\ rate$
@@
Dataset $\downarrow$ $Error\leavevmode\nobreak\ rate$ $Error\leavevmode\nobreak\ rate$ $Error\leavevmode\nobreak\ rate$ $Error\leavevmode\nobreak\ rate$ $Error\leavevmode\nobreak\ rate$ $Error\leavevmode\nobreak\ rate$
@@
Average $Error\leavevmode\nobreak\ rate$ $0.1678$ $0.1775$ $0.1678$ $0.1911$ $0.1552$ $0.1445$
@@
Furthermore, in the analysis of models, we employ the pairwise win-tie-loss (W-T-L) sign test. According to the W-T-L sign test, the null hypothesis posits that two models perform equivalently and are expected to secure victories in $N/2$ datasets, where $N$ denotes the count of datasets. If the classification model secures wins in approximately $N/2+1.96\sqrt{N}/2$ datasets, then the model is deemed significantly better. In the event of an even number of ties between two models, these ties are evenly distributed between them. Nevertheless, if the number of ties is odd, one tie is disregarded, and the remaining ties are distributed among the specified classifiers. In this instance, with $N=55$, the occurrence of wins in at least $42.035$ datasets by one of the models signifies a noteworthy distinction between the models. Table 7 presents the outcomes in terms of pairwise wins, ties, and losses across UCI and KEEL datasets. In each entry, $[x,y,z]$, $x$ denotes the number of times the model specified in the corresponding row secures a win, $y$ represents the number of ties, and $z$ signifies the occurrences of losses in comparison to the model mentioned in the respective column. Upon meticulous examination of Table 7, it is apparent that the proposed MvTPMSVM model outperforms the baseline MvTSVM, MVNPSVM, and PSVM-2V models by achieving a respective number of victories: $47$, $47$, and $43$, out of a total of $55$ datasets. The proposed MvTPMSVM demonstrates a statistically significant distinction compared to the baseline models, except for SVM-2K and MVLDM. Exhibiting a notable level of performance, the MvTPMSVM model succeeds in $39$ out of $55$ datasets over SVM-2K and $26$ out of $55$ datasets over MVLDM. However, the winning percentage of the MvTPMSVM model demonstrates its effectiveness in comparison to the SVM-2K and MVLDM models. It is evident that the proposed MvTPMSVM model exhibits significant superiority when compared to the baseline models.
@@ 4.4 Experiments on real-world AwA datasets
Table 8 displays the ACC, Seny, Spey, and Pren of the proposed MvTPMSVM model along with SVM-2K, MvTSVM, MVNPSVM, PSVM-2V, and MVLDM models. The $error\leavevmode\nobreak\ rates$ and the optimal parameters of the proposed MvTPSVM model and the baseline models are presented in Table 10. From Table 8, it is evident that the proposed MvTPMSVM model exhibits superior generalization performance across the majority of datasets. The average ACC of the proposed MvTPMSVM model and the baseline SVM-2K, MvTSVM, MVNPSVM, PSVM-2V, and MVLDM models are $73.57\%$, $71.02\%$, $69.4\%$, $70.50\%$, $71.02\%$ and $73.33\%$, respectively. Table 8 presents the average rank of all models, determined by their ACC values. It is worth noting that among all the models, our proposed MvTPMSVM model has the lowest average rank. Moreover, we perform the Friedman statistical test, followed by Nemenyi post hoc tests. For the significance level of $\alpha=0.05$, we calculate $\chi_{F}^{2}=50.81$, $F_{F}=12.8345$, and $F_{F}(5,220)=2.2550$. The null hypothesis is rejected as $F_{F}(5,220)<F_{F}$. Now, the Nemenyi post-hoc test is employed to identify significant differences among the pairwise comparisons. We calculate $C.D.=1.1240$, indicating that the average ranking of the models in Table 8 should have a minimum difference by $1.1240$ to be considered statistically significant. The differences in average rank between the proposed MvTPMSVM model and the baseline SVM-2K, MvTSVM, MVNPSVM, PSVM-2V, and MVLDM models are $1.60$, $2.85$, $2.26$, $1.89$ and $0.81$, respectively. $C.D.$ is exceeded by the observed differences. Thus, as per the Nemenyi post hoc test, the proposed MvTPMSVM model, exhibits significant distinctions from SVM-2K, MvTSVM, MVNPSVM, and PSVM-2V except MVLDM. Consequently, the proposed MvTPMSVM model demonstrates superior performance compared to baseline models except MVLDM. However, the proposed MvTPMSVM model surpasses MVLDM in terms of average rank. The evident superiority of the proposed MvTPMSVM model is evident when contrasted with the baseline models.
Table 10: $Error\leavevmode\nobreak\ rate$ and optimal parameters of the proposed MvTPMSVM model along with the baseline models on AwA datasets with the non-linear kernel.
@@
Dataset $\downarrow$ $Error\leavevmode\nobreak\ rate$ $Error\leavevmode\nobreak\ rate$ $Error\leavevmode\nobreak\ rate$ $Error\leavevmode\nobreak\ rate$ $Error\leavevmode\nobreak\ rate$ $Error\leavevmode\nobreak\ rate$
@@
Dataset $\downarrow$ $Error\leavevmode\nobreak\ rate$ $Error\leavevmode\nobreak\ rate$ $Error\leavevmode\nobreak\ rate$ $Error\leavevmode\nobreak\ rate$ $Error\leavevmode\nobreak\ rate$ $Error\leavevmode\nobreak\ rate$
@@
Average $Error\leavevmode\nobreak\ rate$ $0.2808$ $0.3060$ $0.2898$ $0.2950$ $0.2667$ $0.2643$
@@
Subsequently, additional experiments were performed by adjusting the sizes of the training sets, and their respective performance is illustrated in Figures 3. We employ a random selection method, using $30\%$ of the samples for testing, while systematically altering the set sizes allocated for training. Figure 3, exhibiting different training sizes, consistently demonstrates that the proposed MvTPMSVM model generally outperforms the baseline models. Moreover, as the training sizes increase, the MvTPMSVM model exhibits superior performance compared to baseline models.
Based on Figure 5, the Seny, Spey, and Pren analysis indicates that the performance of the proposed MvTPMSVM model achieved the highest position in the baseline models, demonstrating its competitive nature. The average Seny, Spey, and Pren are $79.65$, $78.34$, and $80.36$, respectively. The statistical tests provide further evidence that the proposed models outperformed all the existing models in the AwA datasets and achieved the highest rankings in Seny, Spey, and Pren analyses.